OpenStudy (anonymous):

Which of the following is a solution to the equation log 6 x + log 6 (x + 1) = 1? x = 6/5 x = 2 x = 3 x = 4

OpenStudy (anonymous):

x(x+1)=6^1 x^2+x=6 x^2+x-6=0 Just solve